how much heat would be required to transform 2 kg of saturated liquid-vapor mixture of water at 60% quality and 80 degree celcius to saturated vapour at constant pressure and then to steam at 150 degree celcius and 0.2 MPa at constant volume?


1
Expert's answer
2021-07-08T13:26:19-0400

Gives

M=2kg

"∆m=2\times\frac{60}{100}=1.2kg"

"T_1=80\u00b0c"

"T_2=150\u00b0c"

P=0.2MPa

Q=∆ms ∆T

For steam

"s=1.996KJ\/kgK"


"∆Q=1.2\times1.996(150°-80°)=167.664KJ"